OpenStudy (anonymous):

After 8 months, $750 simple interest was owed ona loan of $9000. Find the annual interest rate.

OpenStudy (gw2011):

You use the following formula: I=(P)(R)(T) Where I=Interest, P=Principal, R=Rate of Interest, T=Time R=I/(P)(T) R=750/(9000)(8/12) R=(750)(12)/(9000)(8) R=9000/72,000 R=0.125=12.5% Therefore, the annual interest is 12.5%